Mr. Agunton is planning for her retirement. He is 30 years old today and would like to have N500,000 when he turns 55, he estimates that he will be able to earn 8 percent rate of return on her retirements over time; he wants to set aside a constant amount of money every year (at the end of the year to help achieve his objective. How much money did Agunfon invest at the end of each of the next 25 years to realize his goal of N500,000 at the end


1
Expert's answer
2021-11-14T17:30:18-0500

"future \space value=R\frac{(1+i)^n-1}{i}\\=500000=R\frac{(1+0.08)^{25}-1}{0.08}\\=500000=R\times 73.10594\\R=6839.3895"